Telescope Performance Verification
with IRTC at LFBG - 15-16 May 2009 MST

Observers:: D.Thompson, J.Heidt (mtn)
Telescope Operator: A. Ceranski
Software Support: C. Biddick (phone)
Telescope Support: T. Shih, J. Little

Summary

Excellent night!
  • Open shutter time 53.4% (see end of this page for details)
  • Presets repeat to ~1 pixel peak-to-peak over 50+ degree EL changes
  • Dice5 patterns behave as expected with rms~45mas

From last night:
  • 1200s guider XY rms

Preparations:

Changed back to good transform (oacontrol.20090515.conf) at Jesper's suggestion.

Preset closed dome (before sunset) ACQUIRE mode to test
18:32:07.671 hangs on OSS "Collimating M3" - restart OSS
18:35:14.007 worked! Ready to go!

Details

18:35:34.028 pointto M5_1246 gs=0 TRACK PAM- PA=0
ptreset to go back to default pointing IE and CA (brought source on-chip)
confirmed with Aaron that all collimation has been cleared.

19:31:49.514 pointto BS9143 Tel="ACQUIRE" gs=0 PAM- PA=0 (too bright to acquire star)
IE changed from -74 to -54, CA from 6.4 to -1

19:42:04.067 pointto BS9143 Tel="ACTIVE" gs=0
19:56:59.883 tweak pointing: IE=-66 CA=-5

Tout=10.1 Tglass=10.0 RH=27.2 some low cirrus at sunset.

20:01:30.810 pointto BS9148 gs=4 EL=78

Restarting IRTC GUI to see if the exptime can be set from iraf...yes. The restarts of iraf must have made the socket connection go stale.

20:10:07.512 pointto BS9148 gs=4 031141 0002_cube 80
20:13:34.798 pointto BS9160 gs=23 M1 hit +Z collimation limit
20:19:36.161 pointto BS9153 gs=8 (mid-EL, to recollimate, pointing OK)

Interleaved presets and dice5 for preset and offset tests

BS9148 starts the night at high EL and BS9169 starts at low EL. We will alternate between presetting and gathering 10x10 dice5 data on these two fields to demonstrate stability of the offsetting and repeatability of the source position on axis as a function of elevation.

20:27:35.110 pointto BS9148 gs=04 033130 0003_cube 84
20:32:48.840 pointto BS9162 gs=01 033715 0004_cube 42
20:38:32.420 pointto BS9169 gs=14 034011 0005_cube 32
20:41:30.032 pointto BS9148 gs=04 034652 0006_cube 86

BS9148 dice5 10x10 30x2.0sec nloop=10 at PA=180.18
  • Start: 034828 0007_cube
  • End: 050001 0056_cube

22:01:16.808 pointto BS9148 gs=04 050202 0057_cube 76
22:03:21.619 pointto BS9169 gs=14 050748 0058_cube 45

offset 4.0 -3.0 coord="RADEC" to center 9169 on-axis star before dice5

BS9169 dice5 10x10 30x2.0sec nloop=10 at PA=316.33
  • Start: 051314 0059_cube
  • End: 062503 0108_cube

23:26:43.882 pointto BS9169 gs=14 062817 0109_cube
23:36:17.464 Correct pointing on pointing start WT10_396 (IE -50, CA -5)
23:39:30.532 pointto BS9169 gs=0 PAM- ACQUIRE (no star?)

23:43:41.371 pointto BS9169 gs=14 064527 0110_cube 58
23:46:52.341 pointto BS9148 gs=04 065012 0111_cube 54

BS9148 dice5 10x10 30x2.0sec nloop=10 at PA=120
  • Start: 065509 0112_cube
  • End: 080506 0161_cube
  • Seeing got progressively worse & variable.
  • In Loop 5 it was often over 2 arcsec on the guider

12:28 MST Tout=8.7 Tprimary=9.5 RH=37.5 Wind 3-9m/s

01:06:23.664 pointto BS9148 gs=04 080844 0162_cube 39
01:10:41.907 pointto BS9169 gs=14 081550 0163_cube 67

Seeing has gotten better, down to ~0.8 on the guider and more stable.

offset 4.0 -3.0 coord="RADEC" to center 9169 on-axis star before dice5

BS9169 dice5 10x10 30x2.0sec nloop=10 at PA=316.33
  • Start: 081801 0164_cube
  • Aborted dice5 at 084208 and 0180_cube as the guide star was lost
  • Strong gradient on guide image and scattered light in WFS

01:46:23.118 pointto WT10_396 ACQUIRE to correct pointing (IE now -70, CA now -7.5)

01:50:41.415 pointto BS9169 gs=14 085331 0181_cube 68
01:54:48.902 pointto BS9148 gs= 4 085801 0182_cube 29

Small-offset dice5 patterns for the rest of the night

01:59:38.372 pointto BS9169 gs=14 PA=330 (off X=0)
BS9169 dice5 0.5x1.5 30x2.0sec nloop=10 at PA=330
  • Start: 090433 0183_cube
  • End: 101620 0232_cube

02:11 Some haze around moon in the east. Not strictly photometric.

BS9169 dice5 0.5x1.5 30x2.0sec nloop=10 at PA=245
  • Start: 102016 0233_cube
  • Aborted: 102147 0234_cube
  • Restart: 103357 0235_cube
  • End: 114711 0284_cube

  • 03:31 an offset looked like it was hanging. We ^C'ed in ira and restarted GCS without tracking down why it appeared to be hung in order to collect enough data here before twilight.

04:48:36.653 pointto BS9169 gs=14 PA=316.33 (default) 115005 0285_cube 52

04:51 Closing dome.

As for last night's overall efficiency, tonight we achieved:
  • 17619 seconds (4.89 hours) of all IRTC data collected tonight
  • 9.15 hours between the sending the first collimation preset to when we closed the dome
  • 4.89/9.15 = 53.4% "open shutter time"

-- DavidThompson - 15 May 2009
Topic revision: r8 - 17 May 2009, DavidThompson
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback